Isolated dc-dc converter is switched at 30 kHz with a duty cycle of 0.65. It is supplying 48 Watts load. The turn’s ratio of the centertapped transformer is 6.5. The dc link of this converter is taken from 48 V battery. The output inductor value is 10 mH. The output inductor current ripple is 12 % of load current. Determine following parameters for half bridge, full bridge and push-pull dc-dc converters (i) output voltage (ii) average and rms current of secondary diode (iii) voltage developed across the primary switch when it is off (iv) peak inverse voltage of secondary diode.


1
Expert's answer
2021-04-28T07:26:21-0400

(i) output voltage

"Z=X_L=2\pi fL=2 \pi \times 30 \times10^3 \times 10 \times 10{^-3}=1884.96 \Omega"

"V_{out}=\frac{1884.96}{1884.96 + 1884.96 \times \frac{12}{100}} \times 48=42.858 V"

(ii) average and RMS current of secondary diode

"I_{Avaerage}= \sqrt{\frac{48}{1884.96}+(\frac{12}{100} \times 0.02546)^2}0.02564 A"

"I_{rms}=0.5 I_{Avaerage}=0.5 \times 0.02564=0.01282 A"

(iii) the voltage developed across the primary switch when it is off

"\frac{V_{PRI}T_{ON}}{N_{PRI}}=\frac{V_{SEC}T_{OFF}}{N_{SEC}}"

"V_{SEC}=\frac{V_{PRI}T_{ON}}{N_{PRI}} \times \frac{V_{SEC}}{T_{OFF}}=48 \times \frac{1}{6.5} \times 0.65 =0.48 V"

(iv) peak inverse voltage of the secondary diode.

"V_{peak}=V_{dc} \pi =48 \pi = 150.796 V"